# Enhancing Urban Visual Place Recognition for Crowdsourced Flood Imagery via LLM-Guided Attention

### v1

@@ Abstract
Crowdsourced street-view imagery from social media provides valuable real-time visual evidence of urban flooding and other crisis events, yet it often lacks reliable geographic metadata for emergency response. Existing Visual Place Recognition (VPR) models exhibit substantial performance degradation when applied to such imagery due to visual distortions and domain shifts inherent in cross-source scenarios. This paper presents VPR-AttLLM, a model-agnostic framework that integrates the semantic reasoning and geospatial knowledge of Large Language Models (LLMs) into established VPR pipelines through attention-guided descriptor enhancement. By leveraging LLMs to identify location-informative regions within the city context and suppress transient visual noise, VPR-AttLLM improves retrieval performance without requiring model retraining or additional data. Comprehensive evaluations are conducted on extended benchmarks including SF-XL enriched with real social-media flood images, synthetic flooding scenarios over established query sets and Mapillary photos, and a new HK-URBAN dataset capturing morphologically distinct cityscapes. Integrating VPR-AttLLM with three state-of-the-art VPR models—CosPlace, EigenPlaces, and SALAD—consistently improves recall performance, yielding relative gains typically between 1–3% and reaching up to 8% on the most challenging real flood imagery. Beyond measurable gains in retrieval accuracy, this study establishes a generalizable paradigm for LLM-guided multimodal fusion in visual retrieval systems. By embedding principles from urban perception theory into attention mechanisms, VPR-AttLLM bridges human-like spatial reasoning with modern VPR architectures. Its plug-and-play design, strong cross-source robustness, and interpretability highlight its potential for scalable urban monitoring and rapid geo-localization of crowdsourced crisis imagery.
@@ 1 Introduction
Street View Imagery (SVI) has become an essential source for urban observation and analysis because it enables detailed, ground-level perspectives on built environments [1, 2]. While platforms such as Google Street View (GSV) provide standardized and extensive spatial coverage, they lack high-frequency temporal updates and cannot capture ephemeral or crisis-related changes in urban scenes—such as flash floods and hurricanes. Consequently, crowdsourced street‑view imagery (SVI) collected through social media and citizen‑reporting platforms has emerged as a valuable complement for timely urban sensing and monitoring [3, 4], enhancing cities’ capacity to assess and respond to rapidly evolving situations, particularly during emergencies [5].
Accurately locating crowdsourced street images is vital for integrating them into spatial decision systems for emergency response and urban resilience planning. This remains challenging because most contributors omit precise geotags for privacy or lack of geospatial literacy for precise locations [6, 7]. Although many cities have deployed official reporting platforms such as New York’s 311 system, persistent reporting friction and limited public awareness constrain participation, leading to biased and incomplete representations of urban issues [8, 9]. During disruptive events such as sudden flooding, citizens often share unstructured photos or videos on social media—data that lack geographic metadata yet offer rich, real‑time visual evidence of urban disruptions (Figure 1). Such imagery complements formal reporting systems by capturing transient, on‑the‑ground scenes that enhance situation awareness. However, current municipal workflows still depend on manual verification, or cross‑referencing by local agencies and journalists, which delays response and limits scalability.
@@
Visual Place Recognition (VPR), also referred to as geo-localization, aims to infer the geographic location of an image by matching it to a geo-tagged reference database. VPR lies at the intersection of computer vision, urban morphology, and environmental perception. Foundational urban perception theories, such as the “Image of the City” [10], "Scale and Psychologies of Space" [11], highlight the role of legibility and structural coherence in spatial recognition, while later computational works have quantified discriminative urban visual elements—demonstrating, for instance, “What makes Paris look like Paris?” [12].
@@
Modern VPR approaches fall broadly into two categories. Retrieval-based methods treat VPR as an image retrieval problem, building city-specific reference databases and retrieving the most similar reference image for a given query—exemplified by NetVLAD [13], CosPlace [14], and more recent SALAD [15]. Classification-based VPR models instead predict a coarse geographic region or city category, allowing global coverage at lower spatial resolution, as demonstrated in PlaNet [16], GeoCLIP [17], and PIGEON [18]. With the progression of deep neural networks, both approaches have achieved significant progress by employing CNN-based backbones (e.g., VGG16, ResNet), vision transformers (e.g., DINO), and multimodal representations (e.g., CLIP).
@@
Despite these advances, robust geo-localization in real-world, crowdsourced SVI remains highly challenging. Even models achieving strong performance on established benchmarks (e.g., SF-XL [14], Pittsburgh-30k [19], Tokyo24/7 [20]) exhibit significant accuracy drops when confronted with distribution shifts, such as extreme weather-induced visual distortions (e.g., flooding, heavy rain) or cross-source variations in data capture devices and environmental conditions. These challenges exemplify a fundamental problem in discriminative AI models: undergeneralization, where model performance degrades considerably when inputs deviate from training distributions [21]. This issue proves particularly acute in VPR for two reasons: (i) extreme weather events generate inherently limited training data due to their rarity, and (ii) the diversity of potential distribution shifts makes scenario-specific retraining computationally prohibitive. Addressing undergeneralization in VPR therefore requires training-free approaches that can dynamically leverage complementary reasoning mechanisms to enhance robustness under challenging conditions without requiring extensive retraining or domain-specific data collection.
@@
The emergence of multimodal Large Language Models (LLMs)—such as GPT-4o, Gemini 2.5, and Qwen-VL—creates new opportunities for resolving these limitations. Beyond text understanding, modern LLMs demonstrate sophisticated scene reasoning and spatial awareness derived from large-scale cross-modal training. They can identify and interpret urban components (e.g., building façades, vegetation, signage, or street furniture) and contextual cues that human observers use to infer place identity. Recent project like GeoIntel [22] show that LLMs can perform coarse geo-reasoning and recognize landmark-rich scenes; however, without integration into established retrieval-based pipelines, such models often lack street-level precision for city-scale VPR tasks.
@@
To address this gap, this paper proposes a VPR enhancement framework with LLM-guided attention (VPR-AttLLM) that leverages LLMs’ semantic reasoning and general geo-knowledge to enhance retrieval-based VPR pipelines under challenging conditions. The central idea is to guide existing pre-trained VPR models with LLM-derived attention map extracted from query images, thereby emphasizing salient regions containing strong location cues (e.g., architectural structure, signage, or horizon geometry) and suppressing noisy regions like flooding street surface (Figure 1). This integration does not require additional training or supporting data, making it model-agnostic and computationally efficient.
@@ 2.1 Visual Place Recognition (VPR) Models
Visual Place Recognition (VPR) is commonly formulated as a large-scale image retrieval task, aiming to match a query image to its corresponding geo-referenced counterpart within a database [23, 20]. The primary challenge is to produce compact, discriminative global descriptors that remain robust to changes in viewpoint, illumination, and environmental conditions [24, 25].
@@
Earlier VPR approaches leveraged handcrafted local descriptors (e.g., SIFT, VLAD), but were later superseded by deep convolutional architectures capable of learning aggregation functions end-to-end. A seminal work, NetVLAD, introduced a differentiable VLAD layer on top of CNN backbones to jointly optimize feature representation and place discrimination [13]. Among subsequent CNN-based methods, a major line of progress centered on learnable pooling functions. The Generalized Mean (GeM) pooling extended conventional max- and average-pooling by learning per-channel exponents, thereby enabling more discriminative global descriptors [26]. Building upon this formulation, classification-based VPR frameworks such as CosPlace and EigenPlaces design geo-balanced training strategies that explicitly exploit the strengths of GeM pooling to improve viewpoint and illumination invariance [14, 27]. These works demonstrate how explicit sampling and hierarchical classification amplify the representation capacity of GeM-aggregated descriptors for robust CNN-based VPR.
@@
A parallel direction emphasizes improving aggregation itself through more expressive clustering or token-based formulations. MixVPR introduced feature mixing for better local–global integration [28], while optimal-transport–based architectures such as SALAD extend this concept by learning score-aware cluster aggregation [15]. Transformer-based models, including DINO-ViT and TransVPR, further increasingly adopted with self-attention mechanisms to capture long-range spatial dependencies and offer token-level feature reasoning [29, 30]. Despite these advances, most current methods remain domain-specialized and exhibit performance degradation under severe environmental or cross-source shifts [31].
@@
Beyond architectural innovations in feature extraction and aggregation, complementary post-processing strategies offer performance improvements without requiring model retraining. Among these approaches, Query Expansion (QE) [32] has emerged as a commonly adopted technique in VPR pipelines. Average Query Expansion (AQE) refines the query descriptor by incorporating weighted averages of top-$k$ retrieved database features, with a tunable parameter $\alpha$ controlling the balance between the original query and retrieved candidates. This simple strategy has demonstrated consistent recall improvements with minimal computational overhead across various VPR benchmarks [33, 34]. However, existing post-processing methods have not been systematically evaluated under challenging cross-domain or extreme weather scenarios with severe appearance variations, and semantic-guided query augmentation strategies that leverage high-level scene understanding remain largely unexplored. Addressing these limitations motivates our proposed VPR-AttLLM framework, which introduces language model guided attention modules for both GeM pooling and cluster-based aggregation, thereby unifying these two architectural paradigms under a single interpretive mechanism. By incorporating semantic reasoning into the feature aggregation process, our approach extends beyond conventional post-processing techniques to enable more adaptive query augmentation across diverse environmental conditions.
@@ 2.2 Integrating LLMs into Street‑View Perception and Geo‑Localization
In recent years, geo-localization has been widely used to evaluate the geographic knowledge and spatial reasoning ability of LLMs, often through GeoGuessr-like experiments. Several studies benchmark LLMs in inferring countries or cities from given street-view imagery using few-shot prompting [35]. For instance, the GeoIntel agent demonstrates LLM-driven inference of global cities [22]. By leveraging textual or visual signatures, Yin et al. [36] achieved over 60% accuracy within 10 km for disaster image localization during Hurricane Harvey.
@@
In parallel, LLMs have demonstrated strong capabilities in interpreting localized urban scenes and performing multimodal classification tasks. They can infer building ages from street-view imagery [37], estimate inundation depth from flooding images [38], or identify improper on-street waste placement [39] through zero- or few-shot reasoning. However, these applications primarily exploit LLMs as semantic decoders that extract or classify local visual attributes within individual images, rather than as spatially contextual components that influence global feature organization. Additionally, although several relevant experiments have been conducted in LLM’s spatial cognition [40] or geo-knowledge [41], existing tests are still constrained in local scenario inference like demographic indicator prediction. Consequently, while LLMs excel at local semantic interpretation, their potential to guide spatial reasoning or retrieval-level representation in VPR pipelines remains unexplored or underutilized.
@@ 3.1 LLM Attention Generation for SVI Query
Building upon this foundational understanding, urban perception theory provides more granular insights into feature distinctiveness. Lynch’s seminal work [10] on urban legibility identifies landmarks—physical references distinguished by their singularity or memorable form—as critical anchors for spatial recognition, with hierarchical distinctiveness: an architecturally unique civic building possesses greater "landmark value" than a generic residential structure, even though both represent permanent infrastructure. Similarly, Golledge’s [42] wayfinding behavior framework and Sorrows and Hirtle’s [43] landmark salience taxonomy, distinguish between perceptually salient elements that serve as distinctive reference points (e.g., ornate building details, unique signage with high visual or cognitive prominence) and background elements forming the spatial context (e.g., repetitive patterns, standard street furniture with low salience). This suggests that finer-grained differentiation based on spatial uniqueness within the broader urban environment—distinguishing common architectural patterns from rare, localizable features—can further enhance place recognition under adverse conditions.
@@
Our framework operationalizes these principles by prompting the LLM to direct attention toward distinctive urban elements rather than generic scene components based on their spatial uniqueness within the broader cityscape context. Rather than simply segmenting objects, the LLM assesses their contribution to place distinctiveness—effectively distinguishing between common architectural patterns abundant citywide and rare, localizable features. LLMs, with their advanced multimodal reasoning capabilities, offer a novel mechanism to provide such context-aware spatial guidance, particularly when domain shifts degrade the reliability of learned visual representations. The complete prompt is provided in the supplementary material B.
@@
A challenge, however, is that although LLMs can analyze images by extracting features from specific regions, they cannot inherently output explicit spatial attention positions. The task of guiding LLMs to generate coordinates with spatial information within images is known as visual grounding and recent research has explored various strategies for aligning LLM responses to pixel-wise regions within images. For instance, GLaMM [44] introduces a unified framework capable of generating natural language responses seamlessly intertwined with pixel-level segmentation masks, enabling grounded conversation generation. Similarly, VPP-LLaVA [45] leverages visual position prompts to enhance LLaVA’s training performance in visual grounding tasks by explicitly encoding spatial cues.
@@
To operationalize this process, we propose an axis-based visual prompting strategy that positions coordinate axes outside the image frame to provide stable spatial references without altering visual content (Figure 2). This design creates a lightweight and interpretable scaffold that enables the LLM’s text-based reasoning to be spatially grounded during inference. The axis-based approach offers clear advantages over alternative visual prompting methods: grid-based overlays introduce visual artifacts that interfere with the SVIs, while bounding-box references, though providing accurate localization, require additional computation and exhibit failure cases in complex scenes. Consistent with insights from prior multimodal grounding research, including visual position prompting methods [45], our approach benefits from explicit positional cues that strengthen the alignment between textual reasoning and visual localization. Qualitative inspection of the LLM-generated attention maps revealed strong alignment between the highlighted regions and the model’s textual reasoning, indicating consistent spatial grounding. These spatial maps encapsulate intermediate spatial reasoning from the LLM and serve as inputs to the subsequent interpolation process for constructing the full attention map.
Given the LLM-generated dictionary containing $N$ spatial attention points, where each point $i$ is defined by a normalized center coordinate $\mathbf{c}_{i}=(x_{i},y_{i})\in[0,1]^{2}$ and an associated attention weight $w_{i}\in[0,2]$, we construct a smooth attention map $\mathbf{A}\in\mathbb{R}^{H\times W}$ over the feature grid using radial basis function (RBF) interpolation. For each spatial location $(h,w)$ in the feature map, we first compute the normalized grid coordinates $\mathbf{g}_{h,w}=(x_{h,w},y_{h,w})$ where $x_{h,w}=\frac{w}{W-1}$ and $y_{h,w}=\frac{h}{H-1}$. The squared Euclidean distance between grid location $(h,w)$ and each center $i$ is computed as $d^{2}_{h,w,i}=\|\mathbf{g}_{h,w}-\mathbf{c}_{i}\|^{2}$. We then apply a Gaussian RBF kernel with bandwidth parameter $\sigma$ to obtain similarity scores:
$s_{h,w,i}=\exp\left(-\frac{d^{2}_{h,w,i}}{2\sigma^{2}}\right)$ (1)
To construct the final attention map, we compute softmax-normalized interpolation weights and aggregate the attention values from all centers:
$a_{h,w,i}=\frac{s_{h,w,i}}{\sum_{j=1}^{N}s_{h,w,j}},\quad\mathbf{A}_{h,w}=\sum_{i=1}^{N}a_{h,w,i}\cdot w_{i}$ (2)
This formulation ensures that locations closer to specified centers receive attention values closer to their associated weights, while distant regions smoothly blend influences from multiple centers. The bandwidth parameter $\sigma=0.2$ was empirically selected through validation experiments to balance localization precision and smooth spatial transitions. Finally, we clamp the resulting attention values to the valid range $[0,2]$ to prevent extrapolation beyond the specified weight bounds. These refined attention maps are subsequently integrated into multiple VPR backbones to evaluate the overall effectiveness of the proposed framework, providing interpretable and spatially grounded attention representations.

### v2

# Enhancing Geo-localization for Crowdsourced Flood Imagery via LLM-Guided Attention

@@ Abstract
Crowdsourced street-view imagery from social media provides valuable real-time visual evidence of urban flooding and other crisis events, yet it often lacks reliable geographic metadata for emergency response. Existing image geo-localization approaches, also known as Visual Place Recognition (VPR) models, exhibit substantial performance degradation when applied to such imagery due to visual distortions and domain shifts inherent in cross-source scenarios. This paper presents VPR-AttLLM, a model-agnostic framework that integrates the semantic reasoning and geospatial knowledge of Large Language Models (LLMs) into established VPR pipelines through attention-guided descriptor enhancement. By leveraging LLMs to identify location-informative regions within the city context and suppress transient visual noise, VPR-AttLLM improves retrieval performance without requiring model retraining or additional data. Comprehensive evaluations are conducted on extended benchmarks including SF-XL enriched with real social-media flood images, synthetic flooding scenarios over established query sets and Mapillary photos, and a new HK-URBAN dataset capturing morphologically distinct cityscapes. Integrating VPR-AttLLM with three state-of-the-art VPR models—CosPlace, EigenPlaces, and SALAD—consistently improves recall performance, yielding relative gains typically between 1–3% and reaching up to 8% on the most challenging real flood imagery. Beyond measurable gains in retrieval accuracy, this study establishes a generalizable paradigm for LLM-guided multimodal fusion in visual retrieval systems. By embedding principles from urban perception theory into attention mechanisms, VPR-AttLLM bridges human-like spatial reasoning with modern VPR architectures. Its plug-and-play design, strong cross-source robustness, and interpretability highlight its potential for scalable urban monitoring and rapid geo-localization of crowdsourced crisis imagery.
@@ 1 Introduction
Street View Imagery (SVI) has become an essential source for urban observation and analysis because it enables detailed, ground-level perspectives on built environments (gebruUsingDeepLearning2017; naikComputerVisionUncovers2017). While platforms such as Google Street View (GSV) provide standardized and extensive spatial coverage, they lack high-frequency temporal updates and cannot capture ephemeral or crisis-related changes in urban scenes—such as flash floods and hurricanes. Consequently, crowdsourced street‑view imagery (SVI) collected through social media and citizen‑reporting platforms has emerged as a valuable complement for timely urban sensing and monitoring (goodchildCitizensSensorsWorld2007; biljeckiStreetViewImagery2021), enhancing cities’ capacity to assess and respond to rapidly evolving situations, particularly during emergencies (qinEnhancingUrbanResilience2025).
Accurately locating crowdsourced street images is vital for integrating them into spatial decision systems for emergency response and urban resilience planning. This remains challenging because most contributors omit precise geotags for privacy or lack of geospatial literacy for precise locations (croitoruGeosocialGaugeSystem2013; tavraRoleCrowdsourcingSocial2021). Although many cities have deployed official reporting platforms such as New York’s 311 system, persistent reporting friction and limited public awareness constrain participation, leading to biased and incomplete representations of urban issues (kontokostaBiasSmartCity2021; boxerEstimatingReportingBias2025). During disruptive events such as sudden flooding, citizens often share unstructured photos or videos on social media—data that lack geographic metadata yet offer rich, real‑time visual evidence of urban disruptions (Figure 1). Such imagery complements formal reporting systems by capturing transient, on‑the‑ground scenes that enhance situation awareness. However, current municipal workflows still depend on manual verification, or cross‑referencing by local agencies and journalists, which delays response and limits scalability. Furthermore, this lack of real-time ground verification disconnects dynamic Urban Digital Twins from the physical reality they aim to simulate, rendering them less effective during rapidly unfolding crises.
@@
Geo-localization, also known as Visual Place Recognition (VPR), aims to infer the geographic location of an image by matching it to a geo-tagged reference database. VPR lies at the intersection of computer vision, urban morphology, and environmental perception. Foundational urban perception theories, such as the “Image of the City” (lynchImageCity2008), "Scale and Psychologies of Space" (montelloScaleMultiplePsychologies1993), highlight the role of legibility and structural coherence in spatial recognition, while later computational works have quantified discriminative urban visual elements—demonstrating, for instance, “What makes Paris look like Paris?” (doerschWhatMakesParis2012).
@@
Modern VPR approaches fall broadly into two categories. Retrieval-based methods treat VPR as an image retrieval problem, building city-specific reference databases and retrieving the most similar reference image for a given query—exemplified by NetVLAD (arandjelovicNetVLADCNNArchitecture2016), CosPlace (bertonRethinkingVisualGeolocalization2022a), and more recent SALAD (izquierdoOptimalTransportAggregation2024). Classification-based VPR models instead predict a coarse geographic region or city category, allowing global coverage at lower spatial resolution, as demonstrated in PlaNet (weyandPlaNetPhotoGeolocation2016), GeoCLIP (cepedaGeoCLIPClipinspiredAlignment2023), and PIGEON (haasPIGEONPredictingImage2024). With the progression of deep neural networks, both approaches have achieved significant progress by employing CNN-based backbones (e.g., VGG16, ResNet), vision transformers (e.g., DINO), and multimodal representations (e.g., CLIP).
@@
Despite these advances, robust geo-localization in real-world, crowdsourced SVI remains highly challenging. Even models achieving strong performance on established benchmarks (e.g., SF-XL (bertonRethinkingVisualGeolocalization2022a), Pittsburgh-30k (toriiVisualPlaceRecognition2013), Tokyo24/7 (Torii-CVPR-2015)) exhibit significant accuracy drops when confronted with distribution shifts, such as extreme weather-induced visual distortions (e.g., flooding, heavy rain) or cross-source variations in data capture devices and environmental conditions. These challenges exemplify a fundamental problem in discriminative AI models: undergeneralization, where model performance degrades considerably when inputs deviate from training distributions (ilievskiAligningGeneralizationHumans2025). This issue proves particularly acute in urban geo-localization for two reasons: (i) extreme weather events generate inherently limited training data due to their rarity, and (ii) the diversity of potential distribution shifts makes scenario-specific retraining computationally prohibitive. Addressing undergeneralization in VPR therefore requires training-free approaches that can dynamically leverage complementary reasoning mechanisms to enhance robustness under challenging conditions without requiring extensive retraining or domain-specific data collection.
@@
The emergence of multimodal Large Language Models (LLMs)—such as GPT-4o, Gemini 2.5, and Qwen-VL—creates new opportunities for resolving these limitations. Beyond text understanding, modern LLMs demonstrate sophisticated scene reasoning and spatial awareness derived from large-scale cross-modal training. They can identify and interpret urban components (e.g., building façades, vegetation, signage, or street furniture) and contextual cues that human observers use to infer place identity. Recent project like GeoIntel (atiilla2024geointel) show that LLMs can perform coarse geo-reasoning and recognize landmark-rich scenes; however, without integration into established retrieval-based pipelines, such models often lack street-level precision for city-scale VPR tasks.
@@
To address this gap, this paper proposes a VPR enhancement framework with LLM-guided attention (VPR-AttLLM). By treating crowdsourced imagery as a sensor network, this framework serves as a robust data assimilation mechanism for Urban Digital Twins, leveraging LLMs’ semantic reasoning and general geo-knowledge to enhance retrieval-based VPR pipelines under challenging conditions. The central idea is to guide existing pre-trained VPR models with LLM-derived attention map extracted from query images, thereby emphasizing salient regions containing strong location cues (e.g., architectural structure, signage, or horizon geometry) and suppressing noisy regions like flooding street surface (Figure 1). This integration does not require additional training or supporting data, making it model-agnostic and computationally efficient.
@@ 2.1 Visual Place Recognition (VPR) Models
Visual Place Recognition (VPR) is commonly formulated as a large-scale image retrieval task, aiming to match a query image to its corresponding geo-referenced counterpart within a database (schindlerCityscaleLocationRecognition2007; torii247Place2018). The primary challenge is to produce compact, discriminative global descriptors that remain robust to changes in viewpoint, illumination, and environmental conditions (zaffarVPRBenchOpenSourceVisual2021; liPlaceRecognitionMeet2025).
@@
Earlier VPR approaches leveraged handcrafted local descriptors (e.g., SIFT, VLAD), but were later superseded by deep convolutional architectures capable of learning aggregation functions end-to-end. A seminal work, NetVLAD, introduced a differentiable VLAD layer on top of CNN backbones to jointly optimize feature representation and place discrimination (arandjelovicNetVLADCNNArchitecture2016). Among subsequent CNN-based methods, a major line of progress centered on learnable pooling functions. The Generalized Mean (GeM) pooling extended conventional max- and average-pooling by learning per-channel exponents, thereby enabling more discriminative global descriptors (radenovicFinetuningCNNImage2018). Building upon this formulation, classification-based VPR frameworks such as CosPlace and EigenPlaces design geo-balanced training strategies that explicitly exploit the strengths of GeM pooling to improve viewpoint and illumination invariance (bertonRethinkingVisualGeolocalization2022a; bertonEigenPlacesTrainingViewpoint2023). These works demonstrate how explicit sampling and hierarchical classification amplify the representation capacity of GeM-aggregated descriptors for robust CNN-based VPR.
@@
A parallel direction emphasizes improving aggregation itself through more expressive clustering or token-based formulations. MixVPR introduced feature mixing for better local–global integration (ali-beyMixVPRFeatureMixing2023), while optimal-transport–based architectures such as SALAD extend this concept by learning score-aware cluster aggregation (izquierdoOptimalTransportAggregation2024). Transformer-based models, including DINO-ViT and TransVPR, further increasingly adopted with self-attention mechanisms to capture long-range spatial dependencies and offer token-level feature reasoning (hauslerPatchNetVLADMultiscaleFusion2021; wangTransVPRTransformerbasedPlace2022). Despite these advances, most current methods remain domain-specialized and exhibit performance degradation under severe environmental or cross-source shifts (gargRevisitAnythingVisual2024).
@@
Beyond architectural innovations in feature extraction and aggregation, complementary post-processing strategies offer performance improvements without requiring model retraining. Among these approaches, Query Expansion (QE) (chumTotalRecallAutomatic2007) has emerged as a commonly adopted technique in VPR pipelines. Average Query Expansion (AQE) refines the query descriptor by incorporating weighted averages of top-$k$ retrieved database features, with a tunable parameter $\alpha$ controlling the balance between the original query and retrieved candidates. This simple strategy has demonstrated consistent recall improvements with minimal computational overhead across various VPR benchmarks (radenovicRevisitingOxfordParis2018; gordoAttentionbasedQueryExpansion2020). However, existing post-processing methods have not been systematically evaluated under challenging cross-domain or extreme weather scenarios with severe appearance variations, and semantic-guided query augmentation strategies that leverage high-level scene understanding remain largely unexplored. Addressing these limitations motivates our proposed VPR-AttLLM framework, which introduces language model guided attention modules for both GeM pooling and cluster-based aggregation, thereby unifying these two architectural paradigms under a single interpretive mechanism. By incorporating semantic reasoning into the feature aggregation process, our approach extends beyond conventional post-processing techniques to enable more adaptive query augmentation across diverse environmental conditions.
@@ 2.2 Integrating LLMs into Street‑View Perception and Geo‑Localization
In recent years, geo-localization has been widely used to evaluate the geographic knowledge and spatial reasoning ability of LLMs, often through GeoGuessr-like experiments. Several studies benchmark LLMs in inferring countries or cities from given street-view imagery using few-shot prompting (wangLLMGeoBenchmarkingLarge2024). For instance, the GeoIntel agent demonstrates LLM-driven inference of global cities (atiilla2024geointel). By leveraging textual or visual signatures, yinLLMenhancedDisasterGeolocalization2025 achieved over 60% accuracy within 10 km for disaster image localization during Hurricane Harvey.
@@
In parallel, LLMs have demonstrated strong capabilities in interpreting localized urban scenes and performing multimodal classification tasks. They can infer building ages from street-view imagery (zengZeroshotBuildingAge2024a), estimate inundation depth from flooding images (lyuAssessingLargeMultimodal2025), or identify improper on-street waste placement (zhangMonitoringStreetlevelImproper2025) through zero- or few-shot reasoning. However, these applications primarily exploit LLMs as semantic decoders that extract or classify local visual attributes within individual images, rather than as spatially contextual components that influence global feature organization. Additionally, although several relevant experiments have been conducted in LLM’s spatial cognition (fengCityGPTEmpoweringUrban2025) or geo-knowledge (manviGeoLLMExtractingGeospatial2024a), existing tests are still constrained in local scenario inference like demographic indicator prediction. Consequently, while LLMs excel at local semantic interpretation, their potential to guide spatial reasoning or retrieval-level representation in VPR pipelines remains unexplored or underutilized.
@@ 3.1 LLM Attention Generation for SVI Query
Building upon this foundational understanding, urban perception theory provides more granular insights into feature distinctiveness. Lynch’s seminal work (lynchImageCity2008) on urban legibility identifies landmarks—physical references distinguished by their singularity or memorable form—as critical anchors for spatial recognition, with hierarchical distinctiveness: an architecturally unique civic building possesses greater "landmark value" than a generic residential structure, even though both represent permanent infrastructure. Similarly, Golledge’s (golledge1999wayfinding) wayfinding behavior framework and Sorrows and Hirtle’s (sorrowsNatureLandmarksReal1999) landmark salience taxonomy, distinguish between perceptually salient elements that serve as distinctive reference points (e.g., ornate building details, unique signage with high visual or cognitive prominence) and background elements forming the spatial context (e.g., repetitive patterns, standard street furniture with low salience). This suggests that finer-grained differentiation based on spatial uniqueness within the broader urban environment—distinguishing common architectural patterns from rare, localizable features—can further enhance place recognition under adverse conditions.
@@
Our framework operationalizes these principles by prompting the LLM to direct attention toward distinctive urban elements rather than generic scene components based on their spatial uniqueness within the broader cityscape context. Rather than simply segmenting objects, the LLM assesses their contribution to place distinctiveness—effectively distinguishing between common architectural patterns abundant citywide and rare, localizable features. LLMs, with their advanced multimodal reasoning capabilities, offer a novel mechanism to provide such context-aware spatial guidance, particularly when domain shifts degrade the reliability of learned visual representations. The complete prompt see Supplementary Material B.
@@
A challenge, however, is that although LLMs can analyze images by extracting features from specific regions, they cannot inherently output explicit spatial attention positions. The task of guiding LLMs to generate coordinates with spatial information within images is known as visual grounding and recent research has explored various strategies for aligning LLM responses to pixel-wise regions within images. For instance, GLaMM (rasheedGLaMMPixelGrounding2024) introduces a unified framework capable of generating natural language responses seamlessly intertwined with pixel-level segmentation masks, enabling grounded conversation generation. Similarly, VPP-LLaVA (tangVisualPositionPrompt2025) leverages visual position prompts to enhance LLaVA’s training performance in visual grounding tasks by explicitly encoding spatial cues.
@@
To operationalize this process, we propose an axis-based visual prompting strategy that positions coordinate axes outside the image frame to provide stable spatial references without altering visual content (Figure 2). This design creates a lightweight and interpretable scaffold that enables the LLM’s text-based reasoning to be spatially grounded during inference. The axis-based approach offers clear advantages over alternative visual prompting methods: grid-based overlays introduce visual artifacts that interfere with the SVIs, while bounding-box references, though providing accurate localization, require additional computation and exhibit failure cases in complex scenes. Consistent with insights from prior multimodal grounding research, including visual position prompting methods (tangVisualPositionPrompt2025), our approach benefits from explicit positional cues that strengthen the alignment between textual reasoning and visual localization. Qualitative inspection of the LLM-generated attention maps revealed strong alignment between the highlighted regions and the model’s textual reasoning, indicating consistent spatial grounding. These spatial maps encapsulate intermediate spatial reasoning from the LLM and serve as inputs to the subsequent interpolation process for constructing the full attention map.
To bridge the gap between semantic reasoning and visual representation, we convert the discrete attention points provided by the LLM into a continuous spatial attention map. The LLM outputs a set of normalized coordinates c_i and associated importance weights w_i. We employ Radial Basis Function (RBF) interpolation with a Gaussian kernel to construct a smooth attention surface $A\in\mathbb{R}^{H\times W}$ over the feature grid.
This process functions similarly to kernel density estimation in spatial analysis: locations explicitly identified by the LLM act as centers of influence, while the RBF kernel ensures a smooth transition of attention values across the image. The bandwidth parameter ($\sigma=0.2$) was empirically selected to balance localization precision with smooth spatial blending. The resulting map is normalized and clamped to the range $[0,2]$, ensuring that the attention mechanism can both up-weight discriminative features and suppress noise without causing numerical instability in the subsequent network layers.